The finite region bounded by the planes z = x, x + z = 8, z = y, y = 8, and z = 0 sketch the region in R3 write the 6 order of integration. No need to evaluate. clear writing please

1
Expert's answer
2022-01-10T16:15:59-0500

The graph of the planes


"z = x, x + z = 8, z = y, y = 8, \text{ and } z = 0"

in "\mathbb{R}^3" is shown below




The order of the integration is shown thus:


"V = \int^{y}_{z=0} \int^{8}_{y=z}\int^{8-z}_{x=z} dxdydz"